Single-asset portfolios: Stocks A, B, and C have expected returns of 15 percent, 15 percent, and 12 percent, respectively, while their standard deviations are 45 percent, 30 percent, and 30 percent, respectively. If you were considering the purchase of each of these stocks as the only holding in your portfolio and the risk-free rate is 0 percent, which stock should you choose?
Answer #1

We would choose the stock with lower value of coefficient of variation because lesser value of coefficient of variation means better risk return trade offs.
Coefficient of variation=(Standard deviation)/(Expected return)

Expected returns and standard deviation for stock A are 15% and 45% respectively
Coefficient of variation=(Standard deviation)/(Expected return)= 45%/15%=3


Expected returns and standard deviation for stock B are 15% and 30% respectively
Coefficient of variation=(Standard deviation)/(Expected return)= 30%/15%=2


Expected returns and standard deviation for stock C are 12% and 30% respectively
Coefficient of variation=(Standard deviation)/(Expected return)= 30%/12%=2.5

As the coefficient of variation is less for stock B, we would choose stock B